# SDEs with supercritical distributional drifts

### v1

@@ 1.1. Main results
known as the Biot-Savart law, and when $\gamma_{j}$ has the same sign, Takanobu [40] established, through a purely probabilistic argument, the existence and uniqueness of a solution that avoids collisions starting from a point $\mathbf{x}=(x_{1},\cdots,x_{N})$ with $x_{i}\neq x_{j}$ (see also [40] for some extensions). For general $\gamma_{j}\in\mathbb{R}$, Osada [31] showed the same result based on heat kernel estimates for generators in a generalized divergence form, obtained in [32], and on potential theoretical results. In our construction, two particles of the solution are allowed to meet, and as a result, the singular point can be touched.
@@ 1.2. Comparison with related works
Subcritical case: $\alpha\in(-\frac{1}{2},0]$. Our result, Theorem 4.8 below, not only encompasses but also extends the work presented in [14, 48] for the range $\alpha\in(-\frac{1}{2},0]$. In particular, when $\alpha=\frac{1}{2}$, our conditions for the coefficient $b$ are not compared to the conditions stated in [48], where $b\in\mathbf{H}^{-\frac{1}{2}}_{p}$ with $p>2d$ is time-independent. Instead, our result allows $b$ to belong to ${\mathbb{L}}^{q}_{T}\mathbf{B}^{-\frac{1}{2}}_{p,1}$ with $\frac{d}{p}+\frac{2}{q}<\frac{1}{2}$, which encompasses cases with time integrability index $q\in(4,\infty]$.
@@
Energy method comparison for $\alpha\in(-1,-\frac{1}{2}]$. In contrast to the energy method employed in [20], our approach differs significantly. We will outline these differences within the context of the following three aspects.
@@
  • Dimensional dependence: Our supercritical condition, which requires $b\in{\mathbb{L}}^{q}_{T}\mathbf{H}^{-1}_{p}$ with $\frac{2}{q}+\frac{d}{p}<1$, is dimension-dependent. This is in contrast to the assumption made in [20, 19], where they require $b\in\mathbf{H}^{-1}_{2+}$, and this assumption is independent of dimension.
  • Initial conditions: In Lemma 5.3 below, our method achieves the tightness without any assumption on the initial data. Furthermore, for the existence of a weak solution and the uniqueness of the generalized martingale solution, we only require that the law of the initial data has an $L^{2}$-density with respect to the Lebesgue measure.
@@
  • Methodology: Our methodologies differ fundamentally from each other. In [20], the focus is on establishing the existence of a unique energy solution on the torus $\mathbb{T}^{d}$. The key aspect of the energy method is to find a stationary distribution $\mu$ and show the following Itô trick:
@@
    ${\mathbb{E}}_{\mu}\left[\sup_{t\in[0,T]}\left|\int_{0}^{t}f(X_{s}){\mathord{{\rm d}}}s\right|^{m}\right]\lesssim_{C}T^{m/2}\|(-{\mathscr{L}}_{s})^{-1/2}f\|_{L^{p}({\mathbb{T}}^{d})}^{p},$
@@
    where ${\mathscr{L}}_{s}:=\frac{1}{2}({\mathscr{L}}+{\mathscr{L}}^{*})$ is the symmetric part of the generator ${\mathscr{L}}:=\Delta+b\cdot\nabla$, which is exactly the Laplacian $\Delta$ when $\mathord{{\rm div}}b=0$. On the other hand, in order to consider the whole space $\mathbb{R}^{d}$, the authors in [19] introduce an additional term $-\frac{1}{n}X_{t}$ to construct the stationary distribution, followed by taking the limit as $n\to\infty$. This approach makes it challenging to address the time-dependent case. In our work, we make use of a classical Krylov-type estimate, where the condition $\frac{2}{q}+\frac{d}{p}<1$ is sharp due to scaling. Reducing the assumption to $p>2$ is difficult within our framework. Importantly, our method does not necessitate the existence of a stationary distribution, which can be challenging in unbounded spaces. Additionally, our approach allows us to handle time-dependent drifts efficiently. Furthermore, our method could be used to deal with the multiplicative noise as well as the degenerate kinetic SDEs, which seems not possible by the energy method developed from [20].
@@
Rough path method. Recently, numerous studies have been dedicated to investigating singular SDEs driven by fractional Brownian motions. Due to the absence of Markov and martingale properties for fractional Brownian motions, traditional stochastic analysis arguments are no longer applicable. In [6, 18], path-by-path well-posedness is established by using the Young integral in rough path theory and the stochastic sewing lemma. It’s important to note that this method only works with subcritical drifts.
@@ 1.3. Organization of the paper
Section 3 is dedicated to the study of the solvability of the linear PDE (1.8) with a distributional drift $b$. In the subcritical case, we utilize Duhamel’s formulation and fixed point theorem, while in the supercritical case, we apply maximum estimates and the energy method.
In Section 4, we focus on proving Theorem 1.2, primarily through the Zvonkin transformation. A crucial step in this process is to establish the following uniform Krylov estimate for the approximation solution $X^{n}$: for any $m\in{\mathbb{N}}$, $0\leqslant t_{0}<t_{1}\leqslant T$ and $f\in{\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}$,
@@
Finally in Section 6, we apply the results to the diffusion in random environments. In particular, vector-valued Gaussian random fields are provided to illustrate our results.
Throughout this paper, we use $C$ with or without subscripts to denote constants, whose values may change from line to line. We also use :=assign:=:= to indicate a definition and set
$a\wedge b:=\max(a,b),\ \ a\vee b:={\mathord{{\rm min}}}(a,b).$
By $A\lesssim_{C}B$ or simply $A\lesssim B$, we mean that for some constant $C\geqslant 1$, $A\leqslant CB$.
2. Prelimiaries
In this section, we recall the definition and some properties of Besov and Bessel potential spaces. Additionally, we establish a Schauder estimate for the heat equation in Besov spaces, which appears to be new. This estimate will be used to solve the Kolmogorov equation in the subcritical case.
@@ 2. Prelimiaries
$\hat{f}(\xi)$ $:=\frac{1}{(2\pi)^{d/2}}\int_{{{\mathbb{R}}^{d}}}\mathrm{e}^{-{\rm i}\xi\cdot x}f(x){\mathord{{\rm d}}}x,\quad\xi\in{{\mathbb{R}}^{d}},$
$\check{f}(x)$ $:=\frac{1}{(2\pi)^{d/2}}\int_{{{\mathbb{R}}^{d}}}\mathrm{e}^{{\rm i}\xi\cdot x}f(\xi){\mathord{{\rm d}}}\xi,\quad x\in{{\mathbb{R}}^{d}}.$
@@
where $(L^{p},\|\cdot\|_{p})$ is the usual $L^{p}$-space in ${\mathbb{R}}^{d}$. Moreover, we use $C_{b}=C_{b}({\mathbb{R}}^{d})$ to denote the Banach space of all bounded continuous functions on ${\mathbb{R}}^{d}$, and $C^{\infty}_{b}=C^{\infty}_{b}({\mathbb{R}}^{d})$ the space of all smooth functions with bounded derivatives of all orders.
@@ 2.1. Besov and Bessel potential spaces
To introduce the Besov space, we first introduce the dyadic partition of unity. Let $\phi_{-1}$ be a symmetric nonnegative $C^{\infty}-$function on $\mathbb{R}^{d}$ with
@@
For given $j\geqslant-1$, the block operator ${\mathcal{R}}_{j}$ is defined on ${\mathscr{S}}^{\prime}({{\mathbb{R}}^{d}})$ by
@@
For $s\in(0,1)$, an equivalent characterization of ${\mathbf{B}}^{s}_{p,q}$ is given by (see [1, p74, Theorem 2.36] or [22, Theorem 2.7])
@@
For $(\alpha,p)\in({\mathbb{R}}\setminus{\mathbb{Z}})\times[1,\infty]$, let ${\mathbf{H}}^{\alpha}_{p}$ be the Bessel potential space defined by:
@@
Note that we do not use the space ${\mathbf{H}}^{\alpha}_{1}$ for $\alpha<0$. Moreover, for integer $\alpha\in{\mathbb{N}}$ and $p\in(1,\infty)$, an equivalent norm in ${\mathbf{H}}^{\alpha}_{p}$ is given by (cf. [37, p135, Theorem 3])
@@
    ${\mathbf{B}}^{0}_{p,1}\hookrightarrow L^{p}\hookrightarrow{\mathbf{B}}^{0}_{p,\infty},\ \ {\mathbf{B}}^{s^{\prime}}_{p,\infty}\hookrightarrow{\mathbf{B}}^{s}_{p,1}\hookrightarrow{\mathbf{B}}^{s}_{p,q}.$ (2.6)
@@
    For $1\leqslant p_{1}\leqslant p\leqslant\infty$, $q\in[1,\infty]$ and $\alpha=\alpha_{1}-\tfrac{d}{p_{1}}+\tfrac{d}{p}$, it holds that (see [5, Theorem 6.5.1])
    $\|f\|_{{\mathbf{B}}^{\alpha}_{p,q}}\lesssim_{C}\|f\|_{{\mathbf{B}}^{\alpha_{1}}_{p_{1},q}},$ (2.8)
@@
    $\|f\|_{{\mathbf{H}}^{\alpha}_{p}}\lesssim_{C}\|f\|_{{\mathbf{H}}^{\alpha_{1}}_{p_{1}}}.$ (2.9)
@@
    $p_{1}\geqslant p,\ p_{2}\geqslant\tfrac{p_{1}}{p_{1}-1},\ \ \tfrac{1}{p}\leqslant\tfrac{1}{p_{1}}+\tfrac{1}{p_{2}}<\tfrac{1}{p}+\tfrac{\alpha}{d},$
    there is a constant $C>0$ such that for all $f\in{\mathbf{H}}^{\alpha}_{p_{1}}$ and $g\in{\mathbf{H}}^{-\alpha}_{p_{2}}$ (see [48, Lemma 2.1]),
    $\|fg\|_{{\mathbf{H}}^{\alpha}_{p}}\lesssim_{C}\|f\|_{{\mathbf{H}}^{\alpha}_{p_{1}}}\|g\|_{{\mathbf{H}}^{-\alpha}_{p_{2}}},$ (2.10)
@@
    $\|f\prec g\|_{{\mathbf{B}}^{\alpha+\beta}_{p,q}}\lesssim_{C}\|f\|_{{\mathbf{B}}^{\beta}_{p_{1},q_{1}}}\|g\|_{{\mathbf{B}}^{\alpha}_{p_{2},q_{2}}}.$
@@
    $\|f\prec g\|_{{\mathbf{B}}^{\alpha}_{p,q}}\lesssim_{C}\|f\|_{p_{1}}\|g\|_{{\mathbf{B}}^{\alpha}_{p_{2},q}}.$
@@
    $\|f\circ g\|_{{\mathbf{B}}^{\alpha+\beta}_{p,q}}\lesssim_{C}\|f\|_{{\mathbf{B}}^{\beta}_{p_{1},q_{1}}}\|g\|_{{\mathbf{B}}^{\alpha}_{p_{2},q_{2}}}.$
@@
    $\|f\circ g\|_{{\mathbf{B}}^{0}_{p,\infty}}\lesssim_{C}\|f\circ g\|_{p}\lesssim_{C}\|f\|_{{\mathbf{B}}^{-\alpha}_{p_{1},q_{1}}}\|g\|_{{\mathbf{B}}^{\alpha}_{p_{1},q_{2}}}.$ (2.14)
@@
For any $\alpha\in(-1,0]$ and $p,q\in[1,\infty]$, there is a $C=C(d,\alpha,p,q)>0$ such that
$\|b\odot\nabla u\|_{{\mathbf{B}}^{\alpha}_{p,q}}\lesssim_{C}\|b\|_{{\mathbf{B}}^{\alpha}_{p,q}}\|u\|_{{\mathbf{B}}^{1+d/p}_{p,1}}$ (2.16)
@@
$\|\mathord{{\rm div}}b\preceq u\|_{{\mathbf{B}}^{0}_{p/2,\infty}}\lesssim_{C}\|\mathord{{\rm div}}b\|_{{\mathbf{B}}^{-\alpha-2}_{p,q/(q-1)}}\|u\|_{{\mathbf{B}}^{2+\alpha}_{p,q}}.$ (2.17)
@@
$\|b\succ\nabla u\|_{{\mathbf{B}}^{\alpha}_{p,q}}\lesssim\|b\|_{{\mathbf{B}}^{\alpha}_{p,q}}\|\nabla u\|_{\infty}\lesssim\|b\|_{{\mathbf{B}}^{\alpha}_{p,q}}\|u\|_{{\mathbf{B}}^{1}_{\infty,1}}.$
Hence, by embedding (2.8),
$\|b\odot\nabla u\|_{{\mathbf{B}}^{\alpha}_{p,q}}\lesssim\|b\|_{{\mathbf{B}}^{\alpha}_{p,q}}\|u\|_{{\mathbf{B}}^{1}_{\infty,1}}\lesssim_{C}\|b\|_{{\mathbf{B}}^{\alpha}_{p,q}}\|u\|_{{\mathbf{B}}^{1+d/p}_{p,1}}.$
For the second estimate (2.17), it follows by Lemma 2.5. ∎

### v3

@@ 1.1. Main results
known as the Biot-Savart law, and when $\gamma_{j}$ have the same sign, Takanobu [42] established, through a purely probabilistic argument, the existence and uniqueness of a solution that avoids collisions starting from a point $\mathbf{x}=(x_{1},\cdots,x_{N})$ with $x_{i}\neq x_{j}$. For general $\gamma_{j}\in\mathbb{R}$, Osada [33] showed the same result based on heat kernel estimates for generators in a generalized divergence form, obtained in [34], and on potential theoretical results. Note that for $i\neq j$ and $M>0$, using Krylov’s estimate (5.10) below with $f({\bf x})=|x_{i}-x_{j}|^{-1/N}{\mathbf{1}}_{|{\bf x}|\leqslant M}$, we obtain for $p>\frac{Nd}{2}$,
$\mathbb{E}\left(\int_{0}^{T}|X^{N,i}_{t}-X^{N,j}_{t}|^{-1/N}{\mathbf{1}}_{\{|\mathbf{X}^{N}_{t}|\leqslant M\}}\,\mathrm{d}t\right)\lesssim\|f\|_{p}<\infty.$
This implies that (×$(\mathbb{P}\times\mathrm{d}t)(\{(\omega,t):X^{N,i}_{t}(\omega)\neq X^{N,j}_{t}(\omega)\})=1$. Unfortunately, within our general framework, it remains unclear whether two particles can meet.
We summarize our main contributions as follows:
  • In Theorem 1.2, we establish the weak well-posedness of SDEs with drift $b=b_{1}+b_{2}$, where $b_{1}\in{\mathbb{L}}^{q_{b}}_{T}{\mathbf{H}}^{\alpha_{b}}_{p_{b}}$ with $\alpha_{b}\in(-1,-\frac{1}{2})$ and $|b_{2}(t,x)|\lesssim 1+|x|$. This result enables us to address cases where $b$ includes Gaussian fields (see Example 6.8 and Subsection 6.3). Potential applications include SDEs in random environments, such as random directed polymers [10], and self-attracting Brownian motion in a random medium [8].
  • In Lemma 5.3, we demonstrate the tightness of the laws of solutions to approximating SDEs with arbitrary initial data and divergence-free drifts in ${\mathbb{L}}^{q_{b}}_{T}{\mathbf{H}}^{-1}_{p_{b}}$, where $2/q_{b}+d/p_{b}<1$. Subsequently, in Theorem 5.10, we prove the well-posedness of the generalized martingale problem for scaling-supercritical SDEs with divergence-free drift and initial data possessing an $L^{2}$ density. As an application, we discuss $N$-particle systems with singular interaction kernels, such as the one described in equation (1.10), which is closely related to the study of propagation of chaos in vortex systems (see [25]).
@@ 1.2. Comparison with related works
Subcritical case: $\alpha\in(-\frac{1}{2},0]$. Our result, Theorem 4.8 below, not only encompasses but also extends the work presented in [13, 50] for the range $\alpha\in(-\frac{1}{2},0]$. In particular, when $\alpha=\frac{1}{2}$, our conditions for the coefficient $b$ are not comparable to the conditions stated in [50], where $b\in\mathbf{H}^{-1/2}_{p}$ with $p>2d$ is time-independent. Instead, our result allows $b$ to belong to ${\mathbb{L}}^{q}_{T}\mathbf{B}^{-1/2}_{p,1}$ with $\frac{d}{p}+\frac{2}{q}<\frac{1}{2}$, which encompasses cases with time integrability index $q\in(4,\infty]$.
@@
Energy method comparison for $\alpha\in(-1,-\frac{1}{2}]$. In contrast to the energy method employed in [19, 20], our approach differs significantly. We will outline these differences within the context of the following three aspects.
  • ${\mathbb{T}}^{d}$ vs ${\mathbb{R}}^{d}$: In this paper, we study the SDE in the full space $\mathbb{R}^{d}$. By contrast, the work in [19] focuses on the $d$-dimensional torus. Similarly, [20] begins with the $d$-dimensional torus and subsequently uses an $m$-periodic torus to approximate $\mathbb{R}^{d}$ as $m\to\infty$. The approach in [19, 20] is motivated by the use of the Lebesgue measure as the initial distribution for the solution. This choice facilitates the application of the forward-backward martingale argument, which is crucial for deriving the Itô trick introduced later in this part.
@@
  • Dimensional dependence: Our supercritical condition, which requires $b\in{\mathbb{L}}^{q}_{T}\mathbf{H}^{-1}_{p}$ with $\frac{2}{q}+\frac{d}{p}<1$, is dimension-dependent. This is in contrast to the assumption made in [19, 20], where they require $b\in{\mathbf{B}}^{-\gamma}_{p,1}$ with $\gamma\in[0,1]$, $\frac{2}{p}\leqslant 1-\gamma$, and this assumption is independent of dimension. It is worth noting that when $b$ is time-independent, the case $b=\mathord{{\rm div}}A$ with some $A\in L^{2}(K)\cap L^{\infty}(K^{C})$, where $K$ is a small compact set, is also studied in [20]. This framework is further extended to the time-dependent case in [18].
@@
  • Methodology: Our methodologies differ fundamentally from each other. In [19, 20], the key aspect of the energy method is to find a stationary distribution $\mu$ and show the following Itô trick:
    ${\mathbb{E}}_{\mu}\left[\sup_{t\in[0,T]}\left|\int_{0}^{t}f(X_{s}){\mathord{{\rm d}}}s\right|^{m}\right]\lesssim T^{m/2}\|(-{\mathscr{L}}_{\rm s})^{-1/2}f\|_{L^{p}({\mathbb{T}}^{d})}^{p},$
    where ${\mathscr{L}}_{\rm s}:=\frac{1}{2}({\mathscr{L}}+{\mathscr{L}}^{*})$ is the symmetric part of the generator ${\mathscr{L}}:=\Delta+b\cdot\nabla$, which is exactly the Laplacian $\Delta$ when $\mathord{{\rm div}}b=0$. In our work, we make use of a classical Krylov-type estimate, where the condition $\frac{2}{q}+\frac{d}{p}<1$ is sharp due to scaling. More precisely, let $p^{\prime},q^{\prime}$ be the conjugate index of $p,q$, and $\varphi_{t}(x)=(2\pi t)^{-d/2}\mathrm{e}^{-|x|^{2}/(2t)}$ be the Gaussian heat kernel. Clearly, we have
@@
    $\left|{\mathbb{E}}\int_{0}^{T}b(t,W_{t}){\mathord{{\rm d}}}t\right|=\left|\int_{0}^{T}\!\!\int_{{\mathbb{R}}^{d}}b(t,x)\varphi_{t}(x){\mathord{{\rm d}}}x{\mathord{{\rm d}}}t\right|\leqslant\|b\|_{{\mathbb{L}}^{q}_{T}{\mathbf{H}}^{-1}_{p}}\|\varphi\|_{{\mathbb{L}}^{q^{\prime}}_{T}{\mathbf{H}}^{1}_{p^{\prime}}}.$
@@
    Since $\|\nabla\varphi_{t}\|_{{p^{\prime}}}=t^{-\frac{1}{2}-\frac{d}{2p}}\|\nabla\varphi_{1}\|_{{p^{\prime}}}$, it is easy to see that $\|\varphi\|_{{\mathbb{L}}^{q^{\prime}}_{T}{\mathbf{H}}^{1}_{p^{\prime}}}$ is finite if and only if $\frac{2}{q^{\prime}}+\frac{d}{p^{\prime}}>d+1$. In other words, the condition $\frac{2}{q}+\frac{d}{p}<1$ is sharp for the above Krylov estimate. Moreover, our method does not necessitate the existence of a stationary distribution, which can be challenging in unbounded spaces. Thus, on the contrary, in Lemma 5.3 below, our method achieves the tightness without any assumption on the initial data. Additionally, our approach allows us to handle time-dependent drifts efficiently. Furthermore, our method could be used to deal with the multiplicative noise as well as the degenerate kinetic SDEs, which seems not possible by the energy method developed from [19, 20].
@@
Rough path method. Recently, numerous studies have been dedicated to investigating singular SDEs driven by fractional Brownian motions. Due to the absence of Markov and martingale properties for fractional Brownian motions, traditional stochastic analysis arguments are no longer applicable. In [17], path-by-path well-posedness is established by using the Young integral in rough path theory and the stochastic sewing lemma. It’s important to note that this method only works with subcritical drifts. For related results on existence and pathwise uniqueness, see also [7].
@@ 1.3. Organization of the paper
Section 3 is dedicated to the study of the solvability of the linear PDE (1.9) with a distributional drift $b$. In the subcritical case, we utilize Duhamel’s formulation and fixed point theorem, while in the supercritical case, we apply maximum principle and the energy method.
In Section 4, we focus on proving Theorem 1.2, primarily through the Zvonkin transformation. A crucial step in this process is to establish the following uniform Krylov estimate for the approximation solution $X^{n}$: for some $m\geqslant 1$ and any $f\in{\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}$ and $0\leqslant t_{0}<t_{1}\leqslant T$,
@@
In Section 6, we apply the results to the diffusion in random environments. In particular, vector-valued Gaussian random fields are provided to illustrate our results.
Finally, in the appendix, we present a decomposition result for a weighted distribution, originally established by Gubinelli and Hofmanovà [21, Lemma 2.4]. For the convenience of readers, we include a detailed proof.
Throughout this paper, we use $C$ with or without subscripts to denote constants, whose values may change from line to line. We also use :=:=:= to indicate a definition and set
$a\vee b:=\max(a,b),\ \ a\wedge b:={\mathord{{\rm min}}}(a,b).$
By $A\lesssim B$, we mean that for some constant $C\geqslant 1$, $A\leqslant CB$.
2. Preliminaries
In this section, we recall the definition and some properties of Besov and Bessel potential spaces. Additionally, we establish a Schauder estimate for the heat equation in Besov spaces. This estimate will be used to solve the Kolmogorov equation in the subcritical case.
@@ 2. Preliminaries
$\hat{f}(\xi):=\frac{1}{(2\pi)^{d/2}}\int_{{{\mathbb{R}}^{d}}}\mathrm{e}^{-{\rm i}\xi\cdot x}f(x){\mathord{{\rm d}}}x,\quad\check{f}(x):=\frac{1}{(2\pi)^{d/2}}\int_{{{\mathbb{R}}^{d}}}\mathrm{e}^{{\rm i}\xi\cdot x}f(\xi){\mathord{{\rm d}}}\xi.$
@@
where $(L^{p},\|\cdot\|_{p})$ is the usual $L^{p}$-space in ${\mathbb{R}}^{d}$. Moreover, we use $C_{b}=C_{b}({\mathbb{R}}^{d})$ to denote the Banach space of all bounded continuous functions on ${\mathbb{R}}^{d}$, and $C^{\infty}_{b}=C^{\infty}_{b}({\mathbb{R}}^{d})$ the space of all smooth functions with bounded derivatives of all orders. Here, we clarify that the notation $\mathbb{L}^{q}_{T}L^{q}=\mathbb{L}^{q}_{T}$ is used throughout the rest of the paper purely for conciseness, making the norm $\|\cdot\|_{\mathbb{L}^{q}_{T}}$ more compact.
@@ 2.1. Besov and Bessel potential spaces
To define Besov spaces, we first introduce dyadic partitions of unity. Let $\phi_{-1}$ be a symmetric nonnegative $C^{\infty}-$function on $\mathbb{R}^{d}$ with
@@
For $j\geqslant-1$, the Littlewood-Paley block operator ${\mathcal{R}}_{j}$ is defined on ${\mathscr{S}}^{\prime}({{\mathbb{R}}^{d}})$ by
@@
For $p=q=\infty$, we simply denote ${\mathbf{C}}^{s}:={\mathbf{B}}^{s}_{\infty,\infty}$.
@@
For $s\in(0,1)$, an equivalent characterization of ${\mathbf{B}}^{s}_{p,q}$ is given by (see [1, p74, Theorem 2.36] or [23, Theorem 2.7])
@@
From this estimate, one sees that ${\mathbf{C}}^{s}$ coincides with the classical Hölder space.
For $(\alpha,p)\in({\mathbb{R}}\setminus{\mathbb{Z}})\times[1,\infty]$, let ${\mathbf{H}}^{\alpha}_{p}$ be the Bessel potential space defined by
@@
Note that we do not use the space ${\mathbf{H}}^{\alpha}_{1}$ for $\alpha<0$. Moreover, for integer $\alpha\in{\mathbb{N}}$ and $p\in(1,\infty)$, an equivalent norm in ${\mathbf{H}}^{\alpha}_{p}$ is given by (cf. [40, p135, Theorem 3])
@@
    ${\mathbf{B}}^{0}_{p,1}\hookrightarrow L^{p}\hookrightarrow{\mathbf{B}}^{0}_{p,\infty},\ \ {\mathbf{B}}^{s^{\prime}}_{p,\infty}\hookrightarrow{\mathbf{B}}^{s}_{p,1}\hookrightarrow{\mathbf{B}}^{s}_{p,q}\hookrightarrow{\mathbf{B}}^{s}_{p,\infty}.$ (2.6)
@@
    For $1\leqslant p_{1}\leqslant p\leqslant\infty$, $q\in[1,\infty]$ and $\alpha\leqslant\alpha_{1}-\tfrac{d}{p_{1}}+\tfrac{d}{p}$, it holds that (see [5, Theorem 6.5.1])
    $\|f\|_{{\mathbf{B}}^{\alpha}_{p,q}}\lesssim\|f\|_{{\mathbf{B}}^{\alpha_{1}}_{p_{1},q}},$ (2.8)
@@
    $\|f\|_{{\mathbf{H}}^{\alpha}_{p}}\lesssim\|f\|_{{\mathbf{H}}^{\alpha_{1}}_{p_{1}}}.$ (2.9)
@@
    $p_{1}\geqslant p,\ p_{2}\geqslant\tfrac{p_{1}}{p_{1}-1},\ \ \tfrac{1}{p}\leqslant\tfrac{1}{p_{1}}+\tfrac{1}{p_{2}}<\tfrac{1}{p}-\tfrac{\alpha}{d},$
    there is a constant $C>0$ such that for all $f\in{\mathbf{H}}^{\alpha}_{p_{1}}$ and $g\in{\mathbf{H}}^{-\alpha}_{p_{2}}$ (see [50, Lemma 2.2]),
    $\|fg\|_{{\mathbf{H}}^{\alpha}_{p}}\leqslant C\|f\|_{{\mathbf{H}}^{\alpha}_{p_{1}}}\|g\|_{{\mathbf{H}}^{-\alpha}_{p_{2}}},$ (2.10)
@@
    $\|f\prec g\|_{{\mathbf{B}}^{\alpha+\beta}_{p,q}}\leqslant C\|f\|_{{\mathbf{B}}^{\beta}_{p_{1},q_{1}}}\|g\|_{{\mathbf{B}}^{\alpha}_{p_{2},q_{2}}}.$
@@
    $\|f\prec g\|_{{\mathbf{B}}^{\alpha}_{p,q}}\leqslant C\|f\|_{p_{1}}\|g\|_{{\mathbf{B}}^{\alpha}_{p_{2},q}}.$
@@
    $\|f\circ g\|_{{\mathbf{B}}^{\alpha+\beta}_{p,q}}\leqslant C\|f\|_{{\mathbf{B}}^{\beta}_{p_{1},q_{1}}}\|g\|_{{\mathbf{B}}^{\alpha}_{p_{2},q_{2}}}.$
@@
    $\|f\circ g\|_{{\mathbf{B}}^{0}_{p,\infty}}\leqslant C\|f\circ g\|_{p}\leqslant C\|f\|_{{\mathbf{B}}^{-\alpha}_{p_{1},q_{1}}}\|g\|_{{\mathbf{B}}^{\alpha}_{p_{1},q_{2}}}.$ (2.14)
@@
We remark that similar decomposition (2.15) has been used in [19, Lemma 2.6].
@@
For $\alpha\in(-1,0]$ and $p,q\in[1,\infty]$, there is a constant $C=C(d,\alpha,p,q)>1$ such that
$\|b\odot\nabla u\|_{{\mathbf{B}}^{\alpha}_{p,q}}\leqslant C\|b\|_{{\mathbf{B}}^{\alpha}_{p,q}}\|u\|_{{\mathbf{H}}^{1}_{\infty}}\leqslant C\|b\|_{{\mathbf{B}}^{\alpha}_{p,q}}\|u\|_{{\mathbf{B}}^{1+d/p}_{p,1}}$ (2.16)
@@
$\|\mathord{{\rm div}}b\preceq u\|_{{\mathbf{B}}^{0}_{p/2,\infty}}\leqslant C\|\mathord{{\rm div}}b\|_{{\mathbf{B}}^{-\alpha-2}_{p,q/(q-1)}}\|u\|_{{\mathbf{B}}^{2+\alpha}_{p,q}}.$ (2.17)
@@
$\|b\succ\nabla u\|_{{\mathbf{B}}^{\alpha}_{p,q}}\lesssim\|b\|_{{\mathbf{B}}^{\alpha}_{p,q}}\|\nabla u\|_{\infty}\lesssim\|b\|_{{\mathbf{B}}^{\alpha}_{p,q}}\|u\|_{{\mathbf{H}}^{1}_{\infty}}.$
Hence, by embedding (2.7) and (2.8),
$\|b\odot\nabla u\|_{{\mathbf{B}}^{\alpha}_{p,q}}\lesssim\|b\|_{{\mathbf{B}}^{\alpha}_{p,q}}\|u\|_{{\mathbf{H}}^{1}_{\infty}}\lesssim\|b\|_{{\mathbf{B}}^{\alpha}_{p,q}}\|u\|_{{\mathbf{B}}^{1}_{\infty,1}}\lesssim\|b\|_{{\mathbf{B}}^{\alpha}_{p,q}}\|u\|_{{\mathbf{B}}^{1+d/p}_{p,1}}.$
For (2.17), by (2.6) and Lemma 2.5-(i) with $p_{1}=p_{2}=p$ and $(q_{1},q_{2})=(q/(q-1),q)$, we have
$\|\mathord{{\rm div}}b\prec u\|_{{\mathbf{B}}^{0}_{p/2,\infty}}\lesssim\|\mathord{{\rm div}}b\prec u\|_{{\mathbf{B}}^{0}_{p/2,1}}\lesssim\|\mathord{{\rm div}}b\|_{{\mathbf{B}}^{-\alpha-2}_{p,q/(q-1)}}\|u\|_{{\mathbf{B}}^{2+\alpha}_{p,q}}.$
For the resonant part, by (2.14) with $p_{1}=p_{2}=p$ and $(q_{1},q_{2})=(q/(q-1),q)$, we have
$\|\mathord{{\rm div}}b\circ u\|_{{\mathbf{B}}^{0}_{p/2,\infty}}\lesssim\|\mathord{{\rm div}}b\|_{{\mathbf{B}}^{-\alpha-2}_{p,q/(q-1)}}\|u\|_{{\mathbf{B}}^{2+\alpha}_{p,q}}.$
Combining the above two estimates, we complete the proof of (2.17). ∎